# “Here’s Your Evidence”: False Consensus in Public Twitter Discussions of COVID-19 Science

### v1

@@ 1 Introduction
As of January 2024, COVID-19 has claimed over 7 million lives worldwide [82]. Despite the pandemic’s severity, misinformation [4, 17, 47] and conspiracy theories [25, 41, 65] have circulated widely, often influencing real-world behavior [41, 66]. Although misinformation can be sourced from untrustworthy individuals or organizations [29, 83], even perceived experts sometimes use academic articles to promote anti-consensus views like opposition to vaccination [37]. This can lead to misunderstandings about scientific support for official guidance since discussions of expert perspectives can be contentious. For instance, after the Great Barrington Declaration [51] suggested shifting away from lockdowns, the John Snow Memorandum countered it, advocating consensus on protective measures against COVID-19 [6].
The pandemic’s heavy politicization [34, 83] further complicates the issue, as attitudes and behaviors differ among partisan groups [12, 39]. Prior beliefs, values, and identities also influence acceptance of misinformation versus evidence-based claims [3, 27], impeding deliberation. E.g., people overwhelmingly target political opponents when correcting misinformation [5], but being corrected leads to elevated political bias in the content that a user shares [56]. Such behavior can hinder effective responses to health crises like COVID-19.
@@
  • RQ2. (a) Are certain sources disproportionately driving public conversations around COVID-19? (b) Is exposure to them associated with whether they espouse unscientific views?
@@
  • RQ4. In what other ways is science related to COVID-19 being (mis)represented in public circles?
@@
Main Findings. Our study yields several main results:
  1. 1.
@@
    New contrarian account sign-ups increase after the pandemic’s onset, with suspicious daily spikes. On Twitter, they sometimes misrepresent or cherry-pick findings from preprints, often against the authors’ advice. Also, they attempt to discredit research opposing contrarian views by suggesting corruption and conflicts of interest.
  4. 4.
@@
    On the other hand, conformist users engage in more constructive discussions, often exploring details or sharing studies revealing nuances in COVID-19 measures.
@@
Contributions. To the best of our knowledge, this mixed-methods study provides the first large-scale examination of COVID-19 science discussion in the public domain, along with an in-depth qualitative analysis of the motivations behind its dissemination. Our work reveals that certain behaviors and patterns are consistent across scientific subjects with established consensus; alas, this consensus is not easily discernible to the general public. In fact, social platforms might often misrepresent it, causing a false impression of scientific division, driven by the disproportionate volume of secondary sources (users, tweets) over primary ones (scientific papers). These misconceptions may normalize fringe viewpoints and bolster anti-institutional sentiment. To counter this, providing accurate consensus estimates to the public, especially during emerging crises like COVID-19, could help reduce confidence in misinformation.
@@ 2.1 False Consensus and its Implications
This may lead to something known as an illusion of consensus; people tend to give uniform weighting to claims with single or multiple primary sources as long as the number of secondary sources citing the primary ones are equal in number [84]. It may occur due to an implicit assumption that the primary sources are not truly independent [23]. Nevertheless, AI fact-checking agents are more effective in their corrections when citing multiple primary sources [75], potentially indicating that illusory consensus is topic-specific or more prominent in human-to-human communications.
Figure 1: Data collection pipeline.
@@ 2.2 Misrepresentation of Science
Total Published Users Retweets Collected in
Papers 25.3k 13.4k N/A N/A Nov 2022
Tweets 1.24M N/A 346k 70.8% Dec 2022
Table 1: Overview of dataset.
Remarks. Scientific inquiry, though focused on accuracy and systematic analysis, may also serve personal sentiments [15, 44, 45] especially when enabled by social environments [3]. This can lead to false perceptions of consensus [23, 27, 26, 38, 84] and influence behavior [41, 66]. This paper quantifies false consensus in COVID-19 discussions to understand if public discourse aligns with scientific evidence. We explore whether sources gain prominence based on the views they support and investigate tactics enabling scientific misinformation. Arguably, our work informs science communication, health policy, and strategies to combat institutional distrust and polarization on scientific matters.
@@ 4.1 Node Classification
Following the preliminary visualization, we treat this as a binary node classification problem. The two possible classes are conformists (i.e., those who follow official guidance) and contrarians (i.e., those who oppose it). We sample 200 random users with $\geq 3$ interactions in the retweet network. Two authors of this paper annotate them based on their tweets’ contents with a Kappa agreement of 0.61, which is considered substantial [48]. Disagreements are resolved through discussion and a final agreement between the two annotators. We do not observe cases where users share or express both contrarian and conformist views during this annotation exercise (e.g., users who are contrarian with respect to vaccines but conformist with respect to masks), suggesting that such cases are exceptionally rare. Thus, we use the full retweet network instead of topic-specific ones for node classification. Moreover, we are able to classify all users as either conformist or contrarian, thus we do not add an “undefined” class.
We extract the network’s giant component and use the adjacency matrix and nodes’ normalized spatial coordinates from a 2D ForceAtlas representation as training features. We then train a 3-layer Graph Convolutional Network with 5-fold validation, where 80% of the data are used for training with the remaining 20% used as a holdout test set at each fold, such that all of the data cycle through both training and testing phases. Weights are reset prior to running every fold so that the model is fully blind to the test data. We achieve an average accuracy of 98% across all folds, suggesting a robust model that is not overfitting to the training data; the final model is trained on all 200 examples.
@@ 5.2 Illusion of Consensus
For each category, we obtain 200 random preprint samples. We then annotate their abstracts on whether they support or oppose official institutions’ standing, relying on category-unique codebooks. Broadly, for each category, we treat official institutions’ standing as 1) supporting the use, efficacy, or safety of vaccines and boosters, 2) supporting the use of interventions and measures to thwart transmission, 3) suggesting that COVID-19 is a potentially dangerous illness, and 4) supporting the use of WHO-approved drugs and opposing the use of non-approved ones. Annotations are performed by two authors, one of whom is a bioengineering expert, and disagreements are resolved through discussion to reach a final consensus (see Table 2 for inter-rater agreements).
Scientific consensus ratios are estimated based on our annotations. These ratios are then matched against the ratios of contrarian to conformist tweets and contrarian to conformist users that tweet about a topic to determine the science-public discrepancy (i.e., false consensus). We additionally annotate the 20 most mentioned papers per category to assess whether their ratios differ from the random samples.
@@
For drug treatments, contrarian discourse on Twitter is more than twice as prominent as expected based on scientific consensus. This category exhibits the lowest discrepancy, possibly because several of the papers we classify as contrarian-consistent are not clinical trials and urge caution before making final recommendations in the drugs they study (see Sections 8.2 and 8.3 for a related discussion). In contrast, vaccine discussions on Twitter exhibit the largest discrepancy, with contrarian tweets being over ten times more prevalent than their consensus-derived ratio. This relative preference for contrarianism remains consistent at the user level.
@@ 5.3 Engagement and Popularity Analyses
[…]
@@ 5.4 Take-Aways
(a) Stacked plot of tweet volume by type.
(b) Unstacked area plot of tweet volume by group.
(c) Stacked plot of volume of paper publications per week.
Figure 7: Time-series plots of tweet volume. Fig. 7(a) includes a comparison to the Altmetric-estimated volume (dotted line).
@@ 7 Tactics and Motivations Behind Science Dissemination
[…]
@@ 7.1 Methodology
In the following, we do not directly quote users to protect their privacy but summarize and describe the most common talking points. For each case, we first overview the paper’s topic and main conclusions and then describe how it is discussed on Twitter.
@@ 7.2 Theme 1: Selective Focus on Subset of Results
Cases 1-4. Though cases in the remainder of this section are examined individually, we discuss cases 1 [36], 2 [13], 3 [24], and 4 [21] for this theme in a “batch,” as they share several common characteristics—more precisely, they are all: 1) conformist-consistent but mostly discussed by contrarians, 2) vaccine-related, and 3) reporting waning vaccine efficacy against variants. Cases 1-3 also report a peculiar “negative effectiveness” whereby vaccinations are associated with increased infection rates several months after immunization, though explicitly warning that this is possibly behavioral or statistical confounding (Case 2 eliminates this finding in a revised version which assesses symptomatic infections specifically.) Case 4 assesses different types of vaccines and finds waning – but still moderate – protection for mRNA vaccines, with very low protection from a Janssen single-dose regiment against variants. Crucially, all four studies report some benefit of vaccines, either against infection or severe disease, especially with booster doses (Cases 1-2). Moreover, they explicitly support continued vaccination and boosting efforts, hence our conformist-consistent annotation.
@@
Case 5. This paper [42] falls under the COVID-19 dangers topic. It is contrarian-consistent but predominantly picked up by conformist users. The paper attempts to estimate relative and absolute COVID-19 mortality risk across several European countries and US states, stratified by age group. It reports that deaths for people under 65 years, especially those with no underlying health conditions, are “remarkably uncommon.” Its first version also reported generally higher relative risk in some US states compared to EU countries for people under 65. However, these figures were updated with new data and more states in a revised version approximately one month later, which lowered relative risk estimates in US states. In both versions, the authors advocate for shielding strategies that focus specifically on protecting the elderly or other high-risk individuals, hence our contrarian-consistent annotation.
@@
Case 6. Also in the COVID-19 dangers topic, this study [7] examines long-COVID prevalence in hospitalized and outpatient COVID-19 cases. It reports a high long-COVID incidence, making it conformist-consistent with respect to COVID-19’s potential danger. However, the study also reports that two vaccine doses predict higher long-COVID odds, which the authors describe as “paradoxical.” Contrarians mostly highlight this, prompting one of the authors to add a comment on the preprint server providing evidence to the contrary and describing public discussions of their work as “unfortunate … cherry-picking”.

### v2

@@ 1 Introduction
As of April 2024, COVID-19 has claimed over 7 million lives worldwide [92]. Despite the pandemic’s severity, misinformation [4, 18, 51] and conspiracy theories [27, 43, 72] have circulated widely, often influencing real-world behavior [43, 73]. Although misinformation can be sourced from untrustworthy individuals or organizations [31, 93], even perceived experts sometimes use academic articles to promote anti-consensus views like opposition to vaccination [39]. This can lead to misunderstandings about scientific support for official guidance since discussions of expert perspectives can be contentious. For instance, after the Great Barrington Declaration [55] suggested shifting away from lockdowns, the John Snow Memorandum countered it, advocating consensus on protective measures against COVID-19 [6].
The pandemic’s heavy politicization [36, 93] further complicates the issue, as attitudes and behaviors differ among partisan groups [13, 41]. Prior beliefs, values, and identities also influence acceptance of misinformation versus evidence-based claims [3, 29], impeding deliberation. For example, people overwhelmingly target political opponents when correcting misinformation [5], but being corrected leads to elevated political bias in the content that a user shares [62]. Such behavior can hinder effective responses to public health crises like COVID-19.
@@
  • RQ2. (a) Are certain sources disproportionately driving COVID-19 public conversations? (b) Is exposure to such sources associated with whether they espouse unscientific views?
@@
  • RQ4. In what other ways has COVID-19 science been (mis)represented in public circles?
@@
Main Findings.
@@
  • New contrarian account sign-ups increase after the pandemic’s onset, with suspicious daily spikes. In Twitter discourse, they sometimes misrepresent or cherry-pick findings from preprints, often against the authors’ advice. Also, they attempt to discredit research opposing contrarian views by suggesting corruption and conflicts of interest.
@@
  • On the other hand, conformist users engage in more constructive scientific discussions, often exploring details or sharing studies revealing nuances in COVID-19 measures.
@@
Contributions. To the best of our knowledge, this mixed-methods study provides the first large-scale examination of COVID-19 science discussion in the public domain, along with an in-depth qualitative analysis of the motivations behind its dissemination. Specifically, we provide the first quantification of COVID-19 scientific consensus, the public understanding of this consensus, and crucially, the juxtaposition between the two. Alas, we find that scientific consensus is misrepresented on social platforms, causing a false impression of scientific division, driven by the disproportionate volume of secondary sources (users, tweets) over primary ones (scientific papers). These misconceptions may normalize fringe viewpoints and bolster anti-institutional sentiment. Moreover, we provide novel accounts of how this misrepresentation is carried out, through an increased volume of new contrarian accounts during the pandemic, as well as a sustained surge in activity of these accounts as the pandemic progresses. Our work also situates established anti-science tactics, such as disingenuous attacks on scientists’ integrity or selective discussion of narrative-affirming science [48], to the COVID-19 context specifically. These findings suggest that providing accurate consensus estimates to the public, especially during emerging crises like COVID-19, could help reduce confidence in misinformation and create more dissonance in movements that paradoxically weaponize science to promote anti-scientific views.
Ethical Statement. This study received ethical approval from the authors’ institution (details are omitted for anonymity purposes). We only collect and use the minimal amount of data required to perform our analyses as per our research questions and follow-up analyses, and avoid direct user quotations to protect their identities.
Disclaimer. Although we use the terms “contrarian” and “conformist” to generally refer to users that are against or for scientific consensus, respectively, we do not wish to imply the ‘correctness’ of their respective views. We only study how science is used holistically among public circles and whether reasonable inferences are drawn from such discussions. For example, it would not be wrong to say that COVID-19 vaccines may lose some efficacy against newer variants. However, it is inaccurate to extrapolate this assertion to claims that COVID-19 vaccines are generally ineffective. “Contrarians” and “conformists” are merely a distinction based on how accurately they portray the state of scientific discourse. We recognize that these terms carry inherent connotations; for example, COVID denialism or anti-establishment ideals may be some of the core narratives motivating the contrarian group. However, we choose these labels due to the high-level nature of our work, as we are not surmizing that the same narratives are necessarily motivating all participants. Assuming that scientific consensus is a reasonable proxy for objective truth with respect to COVID-19, we are instead observing how Twitter discourse represents this truth; that is, whether it contradicts it or conforms to it.
Similarly, we are cautious not to describe scientific papers as contrarian or conformist; rather, we use the terms “contrarian-consistent” and “conformist-consistent.” In this way, we wish to convey that we classify papers based on whether their conclusions align with specific viewpoints and narratives, not on whether papers themselves have a certain “agenda.” Indeed, such a practice would be counter-intuitive. Studying and documenting any risks associated with officially espoused measures, among other things, is crucial to forming robust, evidence-based policy. Challenging conventional wisdom is pivotal in scientific advancement, and labeling science that brings new evidence to bear as “contrarian” would be a misnomer. To this end, we also clarify that we are neither assuming nor implying any authors’ stances or attitudes in classifying preprints. Put simply, we are not raising any issues regarding how COVID-19 science is conducted; this is beyond the scope of our work. We are only studying how this science is disseminated and discussed to ensure that it has the intended impact.
@@ 2.1 False Consensus and its Implications
This may lead to something known as an illusion of consensus; people tend to give uniform weighting to claims with single or multiple primary sources as long as the number of secondary sources citing the primary ones are equal in number [94]. Whereas false consensus describes over-inflated prominence perceptions of certain views, illusory consensus outlines how the perceived credibility of different claims is affected by a fixation on secondary, instead of primary sources in support of them. It may occur due to an implicit assumption that the primary sources are not truly independent [24]. Nevertheless, AI fact-checking agents are more effective in their corrections when citing multiple primary sources [82], potentially indicating that illusory consensus is topic-specific or more prominent in human-to-human communications.
@@ 2.2 Misrepresentation of Science
Figure 1: Data collection pipeline.
Remarks. Scientific inquiry, though focused on accuracy and systematic analysis, may also serve personal sentiments [16, 47, 48] especially when enabled by social environments [3]. This can lead to false perceptions of consensus [24, 29, 28, 40, 94] and influence behavior [43, 73]. This paper is the first effort to quantify false consensus in COVID-19 discussions and the extent to which public COVID-19 science representation is problematic. We explore whether sources gain prominence based on the views they support, as well as how false consensus is propagated through a surge in contrarian activity and contrarian newcomers to the discourse. Furthermore, we investigate tactics enabling scientific misinformation on COVID-19 and uncover that they resemble similar tactics that have been used for decades by traditional anti-science groups, such as the anti-vaccination movement [48]. Arguably, our work informs science communication, health policy, and strategies to combat institutional distrust and polarization on scientific matters.
@@ 3 Dataset
Total Published Users Retweets Collection Date
Papers 25.3k 13.4k N/A N/A 04-Nov-2022
Tweets 1.24M N/A 346k 70.8% 13-Dec-2022
Table 1: Overview of dataset.
@@ 4.1 Node Classification
Following the preliminary visualization, we treat this as a binary node classification problem. The two possible classes are conformists (i.e., those who follow official guidance) and contrarians (i.e., those who oppose it). We sample 200 random users with $\geq 3$ interactions in the retweet network. Two authors annotate them based on their tweets’ contents with a Kappa agreement of 0.61, which is considered substantial [52]. Disagreements are resolved through discussion and a final agreement between the two annotators. We do not observe cases where users share or express both contrarian and conformist views during this annotation exercise (e.g., users who are contrarian with respect to vaccines but conformist with respect to masks), suggesting that such cases are exceptionally rare. Thus, we use the full retweet network instead of topic-specific ones for node classification. Moreover, we are able to classify all users as either conformist or contrarian, thus we do not add an “undefined” class.
We extract the network’s giant component and use the adjacency matrix and nodes’ normalized spatial coordinates from a 2D ForceAtlas representation as training features. We then train a 3-layer Graph Convolutional Network (GCN) with 5-fold validation, where 80% of the data are used for training with the remaining 20% used as a holdout test set at each fold, such that all of the data cycle through both training and testing phases. Weights are reset prior to running every fold so that the model is fully blind to the test data. We achieve an average accuracy of 98% across all folds, suggesting a robust model that is not overfitting to the training data; the final model is trained on all 200 examples.
We elect this user-level network approach instead of a tweet-level language model because the two-community retweet network topology suggests the existence of two contested frames in Twitter discussions of COVID-19 science. Several prior works have successfully utilized retweet networks in such contested frames contexts, like Black Lives Matter [49, 79] or White Helmets discourse [87]. Although these works employ unsupervised, modularity-optimizing algorithms for community detection such as Leiden or Louvain, we use a GCN model for node classification for two reasons. First, this allows us to set the number of potential classes a-priori informed by our own domain expertise, whereas unsupervised community detection methods can result in any number of communities. Second, this approach enables us to manually annotate examples of users for training and testing based on their tweets’ contents. Therefore, despite the actual model relying on the retweet network itself to classify users, the model’s accuracy is based on ground-truth assessments of these users’ tweet contents, which increases our confidence in the model’s performance.
@@ 5.2 Illusion of Consensus
For each category, we obtain 200 random preprint samples. We then annotate their abstracts on whether they support or oppose official institutions’ guidelines at the time of annotation, relying on category-unique codebooks (see auxiliary materials). Broadly, for each category, we treat official institutions’ standing as 1) supporting the use, efficacy, or safety of vaccines and boosters, 2) supporting the use of interventions and measures to thwart transmission, 3) suggesting that COVID-19 is a potentially dangerous illness, and 4) supporting the use of WHO-approved drugs and opposing the use of non-approved ones. Annotations are performed by two authors, one of whom is a bioengineering expert, and disagreements are resolved through discussion to reach a final consensus (see Table 2 for inter-rater agreements).
Scientific consensus ratios are estimated based on our annotations. These ratios are then matched against the ratios of contrarian to conformist tweets (i.e., tweets from contrarian and conformist users, respectively) and contrarian to conformist users that tweet about a topic to determine the science-public discrepancy (i.e., false consensus). We additionally annotate the 20 most mentioned papers per category to assess whether their ratios differ from the random samples.
@@
For drug treatments, contrarian discourse on Twitter is more than twice as prominent as expected based on scientific consensus. This category exhibits the lowest discrepancy, possibly because several of the papers we classify as contrarian-consistent are not clinical trials and urge caution before making final recommendations in the drugs they study (see Section 8.2 for a related discussion). In contrast, vaccine discussions on Twitter exhibit the largest discrepancy, with contrarian tweets being over ten times more prevalent than their consensus-derived ratio. This relative preference for contrarianism remains consistent at the user level.
@@ 5.3 Engagement and Popularity Analyses
Using Altmetric-estimated metrics, we regress the total tweet mentions of each paper on whether the paper is contrarian- or conformist-consistent. We control for the number of authors, news mentions and citations, paper recency, topic, and whether the preprint has been peer-reviewed. We also add citations and Twitter/news mentions from the published version of each paper if one exists. As they are not random with respect to Twitter popularity, we omit the top-20 samples from this analysis. The overall model is significant ($\textit{F}_{(9,612)}=90.81,\textit{p}<0.001,adj.R^{2}=0.239$) and explains 23.9% of the variance. We show all coefficients and confidence estimates in Table 3.
@@
To corroborate this, we also compare contrarian- and conformist-consistent Twitter mentions using a Mann-Whitney U test, which makes no distribution assumptions. This shows that contrarian-consistent papers (median = 193) obtain significantly higher ranks than conformist-consistent ones (median = 28), $\textit{U}=33.0,\textit{p}<0.001$, Common-Language Effect Size (CLES) = 68.5. That is, a randomly-chosen contrarian-consistent paper will accumulate more Twitter mentions than a randomly-chosen conformist-consistent one with a 68.5% probability. However, this advantage does not extend to contrarian users or tweets (see Appendix A). Moreover, we find no evidence that conformist-consistent papers are more likely to pass peer review than contrarian-consistent ones or vice-versa (Appendix B).
@@ 6.1 Tweet and Paper Volume
(a) Stacked plot of tweet volume by type.
(b) Unstacked area plot of tweet volume by group.
(c) Stacked plot of volume of paper publications per week.
Figure 7: Time-series plots of tweet volume. NB: Figure 7(a) includes a comparison to the Altmetric-estimated volume (dotted line).
@@ 7 Tactics and Motivations Behind Science Dissemination
As expected, contrarian- and conformist-consistent papers are mostly shared by contrarians and conformists, respectively. In a Weighted Least Squares (WLS) regression, we confirm that our annotated labels significantly predict paper user distribution (a value from 0 to 1, with 1 meaning a paper is solely shared by contrarians) ($\beta=0.655,\textit{SE}=0.06,\textit{p}<0.001$). The overall model is significant and explains 65% of the variance ($F_{1,622}=116.9,\textit{p}<0.001,R^{2}=0.649$). We weigh the variance contribution of each paper by its number of Twitter mentions.
@@ 7.1 Methodology
In the following, we do not directly quote users to protect their privacy but summarize and describe the most common talking points. For each case, we first overview the paper’s topic and main conclusions and then describe how it is discussed on Twitter. Where cases were published in peer-reviewed journals by the time of data collection, we indicate this along with the publication journal when introducing them. Otherwise, we do not mention publication status.
@@ 7.2 Theme 1: Selective Focus on Subset of Results
Vaccine efficacy against variants. Though cases in the remainder of this section are examined individually, we discuss Cases 1 [38], 2 [14] (later published in JAMA Network Open), 3 [25], and 4 [22] for this theme in a “batch,” as they share several common characteristics—more precisely, they are all: 1) conformist-consistent but mostly discussed by contrarians, 2) vaccine-related, and 3) reporting waning vaccine efficacy against variants. Cases 1-3 also report a peculiar “negative effectiveness” whereby vaccinations are associated with increased infection rates several months after immunization, though explicitly warning that this is possibly behavioral or statistical confounding (Case 2 eliminates this finding in a revised version which assesses symptomatic infections specifically.) Case 4 assesses different types of vaccines and finds waning – but still moderate – protection for mRNA vaccines, with very low protection from a Janssen single-dose regiment against variants. Crucially, all four studies report some benefit of vaccines, either against infection or severe disease, especially with booster doses (Cases 1-2). Moreover, they explicitly support continued vaccination and boosting efforts, hence our conformist-consistent annotation.
@@
Relative vs. absolute mortality. Case 5 [45] (later published in Environmental Research) falls under the COVID-19 dangers topic. It is contrarian-consistent but predominantly picked up by conformist users. The paper attempts to estimate relative and absolute COVID-19 mortality risk across several European countries and US states, stratified by age group. It reports that deaths for people under 65 years, especially those with no underlying health conditions, are “remarkably uncommon.” Its first version also reported generally higher relative risk in some US states compared to EU countries for people under 65. However, these figures were updated with new data and more states in a revised version approximately one month later, which lowered relative risk estimates in US states. In both versions, the authors advocate for shielding strategies that focus specifically on protecting the elderly or other high-risk individuals, hence our contrarian-consistent annotation.
@@
Vaccination and long-COVID odds. Also in the COVID-19 dangers topic, Case 6 [7] examines long-COVID prevalence in hospitalized and outpatient COVID-19 cases. It reports a high long-COVID incidence, making it conformist-consistent with respect to COVID-19’s potential danger. However, the study also reports that two vaccine doses predict higher long-COVID odds, which the authors describe as “paradoxical.” Contrarians mostly highlight this, prompting one of the authors to add a comment on the preprint server providing evidence to the contrary and describing public discussions of their work as “unfortunate … cherry-picking”.